OpenStudy (anonymous):

Solve 4^(x – 2) = 3 using common logarithms.

OpenStudy (anonymous):

\[b^x=A\iff x=\frac{\log(A)}{\log(b)}\]

OpenStudy (anonymous):

\[4^{x-2}=3\iff x-2=\frac{\log(3)}{\log(4)}\]

OpenStudy (anonymous):

"solve using common logs"??!! it makes no difference what log you use

OpenStudy (anonymous):

honestly i dont understand how to do it at all :/

OpenStudy (anonymous):

@satellite73 so wouldnt the log cancel out and it be 3/4??? or is that completely wrong?

OpenStudy (anonymous):

yes, that is completely wrong the log is not a number, it is a function

OpenStudy (anonymous):

you should have the button on your calculator to compute \(\log(3)\)

OpenStudy (anonymous):

so i would do log 3/log 4?

OpenStudy (anonymous):

lets do a simpler example \[2^x=10\] solve in one step using a calculator \[x=\frac{\log(10)}{\log(2)}\]

OpenStudy (anonymous):

3.321928095

OpenStudy (anonymous):

x is the log of the total divided by the log of the base

OpenStudy (anonymous):

so for the original problem would it be....79248125?

OpenStudy (anonymous):

in your case you have \[x-2=\frac{\log(3)}{\log(4)}\] so \[x=\frac{\log(3)}{\log(4)}+2\]

OpenStudy (anonymous):

so 2.792?

OpenStudy (anonymous):

i get this http://www.wolframalpha.com/input/?i=log(3)%2Flog(4)%2B

OpenStudy (anonymous):

alright, thank you so so much :DDD

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

can i ask you one more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

A scientist has 86 grams of a radioactive substance that decays at an exponential rate. Asuming k = –0.4, how many grams of radioactive substance remain after 10 days?

OpenStudy (anonymous):

\[86\times e^{-.4\times 10}=86\times e^{-4}\] and a calculator

OpenStudy (anonymous):

how did you just do that?!

OpenStudy (anonymous):

i am assuming the rate is in days

OpenStudy (anonymous):

yeah, thats what the problem was, i just copied and pasted it.

OpenStudy (anonymous):

\[A_0e^{kt}\] with \(A_0=86\) and \(k=-0.4\)

OpenStudy (anonymous):

and \(t=10\)

OpenStudy (anonymous):

so would it be..195.533

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then what would it be??

OpenStudy (anonymous):

it is decaying so there must be less, not more

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=86e^-4

OpenStudy (anonymous):

so would i do negative 4 then?

OpenStudy (anonymous):

would i subtract that from the 86 then?

OpenStudy (anonymous):

no that is the answer compute it with a calculator

OpenStudy (anonymous):

the answer is \[86\times e^{-4}\] but you need a calculator to compute that number, or use the link i sent

OpenStudy (anonymous):

yeah i got it, thank you so much!

OpenStudy (anonymous):

yw